OpenStudy (kobeni-chan):

Please HELP I've been stuck for an hour ;-; John has taken out a loan for college. He started paying off the loan with a first payment of $100. Each month he pays, he wants to pay back 1.1 times as the amount he paid the month before. Explain to John how to represent his first 20 payments in sigma notation. Then explain how to find the sum of his first 20 payments, using complete sentences. Explain why this series is convergent or divergent.

OpenStudy (nincompoop):

What have you got so far?

OpenStudy (kobeni-chan):

I don't really know where to start :/

OpenStudy (nincompoop):

That is not a good way to solve the problem, dude

OpenStudy (is3535):

100 100*1.1 100*1.1^2 100*1.1^3 ... 100*1.1^19 There they are. You do the sigma notation.

OpenStudy (is3535):

100 + 100*1.1 + 100*1.1^2 + 100*1.1^3 + ... + 100*1.1^19 = 100(1 + 1.1 + 1.1^2 + 1.1^3 + ... + 1.1^19) = \(100\dfrac{1-1.1^{20}}{1-1.1} = 100\cdot\dfrac{5.7275}{0.1} = 5727.50\) You must be familiar with Geometric Series. Very helpful skills.

OpenStudy (is3535):

can you took a hour i took a 3 min

OpenStudy (kobeni-chan):

I'm rusty on a lot of material, I'm not looking for ridicule ;-;

OpenStudy (is3535):

100 + 100*1.1 + 100*1.1^2 + 100*1.1^3 + ... + 100*1.1^19 = 100(1 + 1.1 + 1.1^2 + 1.1^3 + ... + 1.1^19) = 1001−1.1201−1.1=100⋅5.72750.1=5727.50 You must be familiar with Geometric Series. Very helpful skills.

OpenStudy (kobeni-chan):

So I just keep raising the exponent?

OpenStudy (is3535):

57227.50 IS YOUR ANSWER OK MEDAL AND FAN IF YOU WANT

OpenStudy (anonymous):

theres ur answer 57227.50 fan and medal

OpenStudy (kobeni-chan):

I need help getting to the answer! Thats not the point of OpenStudy

OpenStudy (anonymous):

if ur done close the question

OpenStudy (kobeni-chan):

Please, I don't understand though. I get that you guys might be smarter than me or are jsut eager for medals but I need to be able to explain it.

OpenStudy (kobeni-chan):

@ganeshie8 do you know how to explain this?

OpenStudy (is3535):

IM NOT HERE FOR THE MEDAL IM HERE TO HELP

OpenStudy (kobeni-chan):

@is3535 Ok maybe its just your hulk face or something lol

OpenStudy (is3535):

YEA LOL